### v2

@@ 4.5 Illustrative Example
We illustrate Algorithm 1 on a simple example of $p=3$ sensors, where $q=1$ sensor has been corrupted. For the purposes of this illustration in Figure 1, we chose all the sensors to have the same dimension $\mathbb{R}^{n_{y}}$ and the first sensor $y_{1}$ to be corrupted. For each sensor $i$, we obtain its corresponding output measurement set $\mathcal{Y}_{k}^{i}$ in the output space $\mathbb{R}^{n_{y}}$ as depicted in Figure 1, which do not necessarily intersect as each sensor $i$ can be measuring different components of the state $x(k)\in\mathbb{R}^{n_{x}}$.
@@
Next, the time update set $\hat{\mathcal{X}}_{k|k-1}$ is constructed according to (5) and shown in the second box from the top in Figure 1. Each output measurement set $\mathcal{Y}_{k}^{i}$ is mapped back to the state space $\mathbb{R}^{n_{x}}$ and constrained to be in the time update set $\hat{\mathcal{X}}_{k|k-1}$ as shown in the third box from the top in Figure 1. The intersection of the measurement update sets $\hat{\mathcal{Z}}^{i}_{k}$ is informative as the intersections of the attack-free $\hat{\mathcal{Z}}^{i}_{k}$ is non-empty. However, the measurement update set $\hat{\mathcal{Z}}^{i}_{k}$ corresponding to a sensor $i$ that is corrupted could also intersect with the other attack-free measurement update sets. This scenario is depicted in Figure 1 where the measurement update set $\hat{\mathcal{Z}}^{1}_{k}$ of the corrupted sensor $1$ intersects with the attack-free measurement update set $\hat{\mathcal{Z}}^{2}_{k}$ due to an intelligently designed attack signal $a_{1}$. We discuss classes of attack signals which can be detected by our algorithm in Section 5.
@@
Since the standing assumption of the algorithm is that the identity of the corrupted sensor(s) is unknown at any given time step $k$, but the number of corrupted sensors $q$ is known, we have to check for non-empty intersections of all $p-q$ combinations of the measurement update sets $\hat{\mathcal{Z}}_{k}^{i}$, which totals to $\binom{p}{p-q}$ checks. Its union then forms our estimated set $\hat{\mathcal{X}}_{k}$ as depicted in the bottom box of Figure 1.
@@
Figure 1: Illustration of Algorithm 1 with $p=3$ and $q=1$. The algorithm deduces that either sensor $1$ or $3$ has been compromised since their corresponding measurement update sets $\hat{\mathcal{Z}}^{1}_{k}$ and $\hat{\mathcal{Z}}^{3}_{k}$ do not intersect.
4.6 Advantage of S3E Algorithm over Point-based Methods
@@
$\|\hat{x}(k)-x(k)\|\leq\beta(\|\hat{x}(0)-x(0)\|,k)+\gamma_{1}(\|w\|_{\ell^{\infty}})+\gamma_{2}(\|v\|_{\ell^{\infty}})+\gamma_{3}(\|a\|_{\ell^{\infty}})\mathbb{1}_{q\geq\frac{p}{2}}$ (10)
@@
where $\beta$ is a $\mathcal{KL}$ function, $\gamma_{1},\gamma_{2},\gamma_{3}$ are $\mathcal{K}_{\infty}$ functions [^3] A continuous function $\gamma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, if it is strictly increasing and $\gamma(0)=0$. It is a class $\mathcal{K}_{\infty}$ function if it is also unbounded. A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{KL}$ function, if: (i) $\beta(.,s)$ is a class $\mathcal{K}$ function for any $s\geq 0$; (ii) $\beta(r,.)$ is non-increasing and (iii) $\beta(r,s)\to 0$ as $s\to\infty$ for any $r\geq 0$., and $\mathbb{1}_{q\geq\frac{p}{2}}$ is the indicator function, which is $1$ when $q\geq\frac{p}{2}$ and 00 otherwise. Here, $v=[\begin{array}[]{ccc}v_{1}^{\top}&\dots&v_{p}^{\top}\end{array}]^{\top}$ and $a=[\begin{array}[]{ccc}a_{1}^{\top}&\dots&a_{p}^{\top}\end{array}]^{\top}$ concatenate the measurement noise and attack signals in all sensors in a column vector, respectively.
@@
It is easy to see that if the number of sensors $p$ is not strictly twice the number of corrupted sensors $q$ (i.e., the assumption of $q<p/2$ is violated), the estimation error bound depends on the magnitude of the attack signal. In this case, the attacker can arbitrarily destroy the accuracy of point-based estimators by injecting attack signals of very large magnitude. Our set-based approach, on the other hand, avoids this situation because it restricts the attack space by only allowing signals of small magnitude. Large attack signals are automatically discarded by either the measurement update (7) or the agreement protocol (8).
@@
Another important advantage of our set-based approach is that the estimated set $\hat{\mathcal{X}}_{k}$ contains a collection of constrained zonotopes, which serve as multiple hypotheses for the true state $x(k)$ in the state space. Although beyond the scope of this paper, an interesting research direction is to developing a hypothesis testing technique to figure out the true set from the sets influenced by the attack signals. For point-based estimators, on the other, the estimation error bound (10) provides a single box of where the state might be. Moreover, the bounds like (10) that are based on comparison functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ are good for theoretical analysis but are usually very conservative in practice. Therefore, even under no attacks, the value of the right-hand side of (10) is very large as compared to the set based bound presented in the previous section.
@@ 4.7 Bound on the Set-based Estimation Error
Since Theorem 2 guarantees that the true state $x(k)$ of system (3) lies in at least one of the constrained zonotopes in the estimated collection of sets $\hat{\mathcal{X}}_{k}$ at each $k\in\mathbb{Z}_{\geq 0}$, it must also lie in a zonotope that overbounds $\hat{\mathcal{X}}_{k}$. That is, let $\hat{\mathcal{X}}_{k}$ be overbounded by a zonotope
@@
Then, the estimation error can be bounded by
@@
The stability analysis of this error bound is provided in Section 6 by using a point-based secure estimator. Nonetheless, we remark that this error bound is significantly smaller in practice than the error bounds obtained by point-based secure estimators.
4.8 Methods to Improve Computational Complexity
A major drawback of Algorithm 1 is that the attacker can design intelligent and stealthy attacks such that the number of constrained zonotopes in the estimate $\hat{\mathcal{X}}_{k}$ keep growing with time. Although generating such attacks is a difficult problem as we argue in Section 5, the computational complexity of the algorithm may increase exponentially with time $k$ in the worst case scenario. Therefore, it is important to reduce the complexity by adopting several pruning methods that are described below.
The first obvious step is to remove any empty sets or subsets of other sets in the estimated set $\hat{\mathcal{X}}_{k}$. It is also possible to obtain a single overbounding zonotope [4] of $\hat{\mathcal{X}}_{k}$ as in (12), and use it instead of $\hat{\mathcal{X}}_{k}$ in the next time update. However, a better trade-off between accuracy and complexity is to not overbound the whole collection but only the intersecting zonotopes in the collection $\hat{\mathcal{X}}_{k}$. This may not make the cardinality of $\hat{\mathcal{X}}_{k}$ equal to one, but it reduces it significantly by allowing minimal loss of accuracy.
Another method is to employ zonotope reduction methods [40] to reduce the number of generators in the zonotopes, which are often increased when performing the Minkowski sum operations. However, this technique may result in a larger radius of $\hat{\mathcal{X}}_{k}$ and can also yield conservative estimates.
Finally, one could employ a secure point-based state estimator, if it exists, in parallel with the set-based secure estimator. In this case, we may consider only those candidates in the measurement update collection that lie within the intersection of $\hat{\mathcal{X}}_{k}$ and an error margin generated by a secure point-based state estimator. However, as argued before, the existing point-based estimators require that the total number of sensors be strictly greater than twice the number of compromised sensors $2q<p$ and also some works assume that the members of the attacked sensors also remain unchanged over time, which are tighter requirements than our standing Assumption 1(i). Moreover, the error margins obtained by point-based estimators are usually very conservative. Nonetheless, point-based techniques are useful in the stability analysis of the proposed secure set-based state estimator, as discussed in Section 6.
@@ 5 Attack Detection, Identification, and Filtering
In this section, we provide sufficient conditions that guarantee security of the S3E algorithm. First, we provide a condition under which we can detect whether there is an attack on at least one sensor. Second, we provide lower bounds on the magnitude of the attack signals, which, if satisfied, allow us to identify the attacked sensor and/or filter the attack signal to render it ineffective. An attack can be detected whenever the estimated set contains more than one non-intersecting constrained zonotopes. However, an attack is only identified and filtered out either in the measurement update or in the agreement protocol. We also provide a simple attack identification algorithm based on the derived sufficient conditions. The attacks that do not satisfy these conditions are considered to be stealthy, which may be undetectable under the current framework.
@@ 5.1 Attack Detection
The above proposition shows that whenever the estimated set contains more than one constrained zonotopes that are disjoint, there has to be at least one sensor that is compromised. Once the attack is detected, the system operator can take necessary measures to secure the system and bring it to safety. If the attacker wants to stay stealthy, they must design attack signals that ensure that all the agreement sets yield a non-empty intersection. In the next subsections, we show that the magnitude of such a stealthy attack has to be sufficiently small.
@@ 5.2 Attack Identification
We say that the attack $a_{i}(k)$ is identified if the system operator knows for certain that sensor $i$ is compromised at time $k$, i.e., $i\notin\mathsf{S}_{k}$. The attack identification allows us to automatically discard the measurements of the compromised sensors from the S3E algorithm. We achieve attack identification through a measurement update (7), which involves a generalized intersection between the time update $\hat{\mathcal{X}}_{k|k-1}$ and the output measurement set $\mathcal{Y}_{k}^{i}$.
@@
Given a zonotope $C_{i}\bar{\mathcal{X}}_{k|k-1}$ with center $C_{i}c_{x}(k|k-1)$ and radius $\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$, we have that any point $p\in\mathbb{R}^{m_{i}}$ is outside the $C_{i}\bar{\mathcal{X}}_{k|k-1}$ if $\|C_{i}c_{x}(k|k-1)-p\|>\sqrt{m_{i}}\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$. Similarly, for $\mathcal{Y}_{k}^{i}$ whose center is $C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)$ and radius is $\mathrm{rad}(\mathcal{Y}_{k}^{i})$, any $p\in\mathbb{R}^{m_{i}}$ is outside $\mathcal{Y}_{k}^{i}$ if $\|C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)-p\|>\sqrt{m_{i}}\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Here, the square root term $\sqrt{m_{i}}$ appears because the radius of a zonotope is a maximum norm, which relates to the Euclidean norm as follows: for any vector $p\in\mathbb{R}^{m_{i}}$, $\|p\|\leq\sqrt{m_{i}}\|p\|_{\infty}$.
If the distance between the centers of zonotopes $C_{i}\bar{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ is more than $\sqrt{m_{i}}[\mathrm{rad}(\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{Y}_{k}^{i})]$, then any point $x\in\bar{\mathcal{X}}_{k|k-1}$ will be such that $C_{i}x\notin\mathcal{Y}_{k}^{i}$. Therefore, in this case, the measurement update (7) is empty. In other words, we have $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$ if
@@
The sufficient condition in Theorem 3 provides a lower bound for the attack signals that can be identified through a measurement update. If the measurement update set $\hat{\mathcal{Z}}_{k}^{i}$ is empty, then it reveals that sensor $i$ is compromised at time $k$. The implication does not hold in the other direction.
If the attacker wants to stay stealthy and effective, it is necessary to violate the lower bound (14) by choosing attack signals of smaller magnitude. If (14) holds, then the attack $a_{i}(k)$ will be automatically identified and filtered out at the measurement update by yielding $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$.
Moreover, notice that the bound (14) depends on the realization of measurement noise $v_{i}(k)$, which is unknown to both the attacker and the defender. One can lower bound (14) by using $\mathrm{rad}(\mathcal{V}_{i})$ and obtain a deterministic bound, which the attacker can violate to generate stealthy attack signals, as follows
@@
Thus, if the attacker would like to stay stealthy by avoiding an empty measurement update set, they must generate an attack $a_{i}(k)$ that satisfies
@@
Finally, notice that satisfying (14) is a sufficient condition for an empty $\hat{\mathcal{Z}}_{k}^{i}$. Therefore, from the attacker’s perspective, violating (14) is a necessary condition for non-empty $\hat{\mathcal{Z}}_{k}^{i}$. This means that if the attack signal violates (14), the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ could still be empty if $C_{i}\hat{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ are inconsistent with each other at time $k$.

### v3

@@ 4.5 Illustrative and numerical examples
We illustrate Algorithm 1 on a simple example of $p=3$ sensors, where $q=1$ sensor has been compromised. For the purposes of this illustration in Fig. 1, we chose all the sensors to have the same dimension $\mathbb{R}^{n_{y}}$ and the first sensor to be compromised. For each sensor $i$, we obtain its corresponding output measurement set $\mathcal{Y}_{k}^{i}$ in the output space as depicted in Fig. 1, which do not necessarily intersect as each sensor $i$ can be measuring different components of the state $x(k)\in\mathbb{R}^{n_{x}}$.
@@
Next, the time update set $\hat{\mathcal{X}}_{k|k-1}$ is constructed according to (4) and shown in the second box from the top in Fig. 1. Each output measurement set $\mathcal{Y}_{k}^{i}$ is mapped back to the state space $\mathbb{R}^{n_{x}}$ and constrained to be in the time update set $\hat{\mathcal{X}}_{k|k-1}$ as shown in the third box from the top in Fig. 1. The intersection of the measurement update sets $\hat{\mathcal{Z}}^{i}_{k}$ is informative as the intersections of the attack-free $\hat{\mathcal{Z}}^{i}_{k}$ is non-empty. However, the measurement update set $\hat{\mathcal{Z}}^{i}_{k}$ corresponding to a sensor $i$ that is corrupted could also intersect with the other attack-free measurement update sets. This scenario is depicted in Fig. 1 where the measurement update set $\hat{\mathcal{Z}}^{1}_{k}$ of the compromised sensor $1$ intersects with the attack-free measurement update set $\hat{\mathcal{Z}}^{2}_{k}$ due to an intelligently designed attack signal $a_{1}$. We discuss classes of attack signals that can be detected by our algorithm in Section 5.
@@
Since the identity of the compromised sensor(s) is unknown at any given time step $k$, but the number of compromised sensors $q$ is known, we have to check for non-empty intersections of all pqp-q combinations of the measurement update sets $\hat{\mathcal{Z}}_{k}^{i}$, which totals to $\binom{p}{p-q}$ checks. Its union then forms our estimated set $\hat{\mathcal{X}}_{k}$ as depicted in the bottom box of Fig. 1.
@@
Figure 1: Illustration of Algorithm 1 with $p=3$ and $q=1$. The algorithm deduces that either sensor 11 or 33 has been compromised since their corresponding measurement update sets $\hat{\mathcal{Z}}^{1}_{k}$ and $\hat{\mathcal{Z}}^{3}_{k}$ do not intersect.
To further elucidate Algorithm 1, we consider a simple numerical example. Simulations for more realistic systems are provided later in Section 6.
Example 1.
Consider the following system
$x(k+1)$ $=$ $0.1x(k)+w(k),$
$y_{i}(k)$ $=$ $x(k)+v_{i}(k)+a_{i}(k),\qquad i\in\{1,2,3\}$
where for all $k\in\mathbb{Z}_{\geq 0}$, the disturbance and noise are $w(k)\in[-10,10]$ and $v_{i}(k)\in[-0.1,0.1]$, respectively. Notice that this system is observable via each sensor $i$. For the purposes of this illustration, suppose $x(0)=0$, $\hat{\mathcal{X}}_{0}=[-100,100]$ and the disturbance and noise at time step $k=1$ are $w(1)=4$, $v_{1}(1)=0.03$, $v_{2}(1)=-0.02$, $v_{3}(1)=-0.07$, respectively. Notice also that $x(1)=4$.
Suppose sensor $1$ and $2$ are under attack and sensor $3$ is uncompromised, i.e., $p=3$, $q=2$, $c_{\mathsf{J}}=1$, the parameter of Algorithm 1 is $n_{\mathsf{J}}=3$, and $a_{3}(k)=0$ for all $k\in\mathbb{Z}_{\geq 0}$. Further, suppose the attacks on sensors $1$ and 22 at time step $k=1$ are $a_{1}(1)=-10.03$ and $a_{2}(1)=-10$. We construct the following sets according to Algorithm 1 for time step $k=1$:
  • The measurement output set for each sensor according to (5) is $\mathcal{Y}^{1}_{k}=[-6.1,-5.9]$, $\mathcal{Y}^{2}_{k}=[-6.12,-5.92]$ and $\mathcal{Y}^{3}_{k}=[3.83,4.03]$, respectively.
  • The time update set according to (4): $\hat{\mathcal{X}}_{1|0}=0.1\hat{\mathcal{X}}_{0}\oplus\mathcal{W}=[-10,10]\oplus[-10,10]=[-20,20]$.
  • Measurement update sets according to (6): $\hat{\mathcal{Z}}^{1}_{1}=[-6.1,-5.9]$, $\hat{\mathcal{Z}}^{2}_{1}=[-6.12,-5.92]$, $\hat{\mathcal{Z}}^{3}_{1}=[3.83,4.03]$.
  • Agreement sets using (7): for $\mathsf{J}_{1}=\{1\}$, $\mathsf{J}_{2}=\{2\}$ and $\mathsf{J}_{3}=\{3\}$, we get $\mathcal{I}^{1}_{1}=\hat{\mathcal{Z}}_{1}=[-6.1,-5.9]$, $\mathcal{I}^{2}_{1}=\hat{\mathcal{Z}}^{2}_{1}=[-6.12,-5.92]$, and $\mathcal{I}^{3}_{1}=\hat{\mathcal{Z}}^{3}_{1}=[3.83,4.03]$.
  • Finally, the estimated set using (8) is $\hat{\mathcal{X}}_{1}=\mathcal{I}^{1}_{1}\cup\mathcal{I}^{2}_{1}\cup\mathcal{I}^{3}_{1}=[-6.12,-5.9]\cup[3.83,4.03]$.
By Theorem 2, the true state $x(1)$ at time step $k=1$ can lie either in the interval [6.12,5.9][-6.12,-5.9] or $[3.83,4.03]$. The disparity between the two estimated intervals highlights the limitation of the S3E algorithm: is the true state contained in one interval or the other? Nonetheless, having multiple sets of reasonable sizes that can potentially contain the true state provide utility in safety-critical applications, where the estimated set serves to verify whether the system’s trajectory could have entered regions of the state space that are designated as unsafe.
4.6 S3E Algorithm vs. Point-based Methods
@@
$\|\hat{x}(k)-x(k)\|\leq\beta(\|\hat{x}(0)-x(0)\|,k)+\gamma_{1}(\|w\|_{\ell^{\infty}})+\gamma_{2}(\|v\|_{\ell^{\infty}})+\gamma_{3}(\|a\|_{\ell^{\infty}})\mathbbm{1}_{q\geq\frac{p}{2}}$ (9)
@@
where $\beta$ is a $\mathcal{KL}$ function, $\gamma_{1},\gamma_{2},\gamma_{3}$ are $\mathcal{K}_{\infty}$ functions [^2] A continuous function $\gamma:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{K}$ function, if it is strictly increasing and $\gamma(0)=0$. It is a class $\mathcal{K}_{\infty}$ function if it is also unbounded. A continuous function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is a class $\mathcal{KL}$ function, if: (i) $\beta(.,s)$ is a class $\mathcal{K}$ function for any $s\geq 0$; (ii) $\beta(r,.)$ is non-increasing and (iii) $\beta(r,s)\to 0$ as $s\to\infty$ for any $r\geq 0$., and $\mathbbm{1}_{q\geq\frac{p}{2}}$ is the indicator function, which is $1$ when $q\geq\frac{p}{2}$ and 0 otherwise. Here, $v=[\begin{array}[]{ccc}v_{1}^{\top}&\dots&v_{p}^{\top}\end{array}]^{\top}$ and $a=[\begin{array}[]{ccc}a_{1}^{\top}&\dots&a_{p}^{\top}\end{array}]^{\top}$ concatenate the measurement noise and attack signals in a column vector, respectively.
@@
It is easy to see that if the number of sensors $p$ is not strictly twice the number of compromised sensors $q$ (i.e., the assumption of $q<p/2$ is violated), the estimation error bound depends on the magnitude of the attack signal. In this case, the attacker can arbitrarily degrade the accuracy of point-based estimators by injecting attack signals of arbitrarily large magnitude. On the other hand, our set-based approach avoids this situation because it restricts the attack space by allowing only the signals of small magnitude to go undetected. Large attack signals are automatically detected and discarded by either the measurement update (6) or the agreement protocol (7).
@@
Another important advantage of our set-based approach is that the estimated set $\hat{\mathcal{X}}_{k}$ contains a collection of agreement sets, which serve as multiple hypotheses for the true state $x(k)$ in the state space [^3] Although beyond the scope of this paper, an interesting research direction is developing a hypothesis testing technique to figure out the true set from the sets influenced by the attack signals.. Point-based estimators, on the other hand, have an estimation error bound (9) that indicates a single ball inside which the true state might lie. In the presence of attacks on more than half the number of sensors, the size of this ball depends on the magnitude of the attacks. Thus, as described in Example 1, our set-based method can at least identify plausible regions in the state space containing the true state. Moreover, even when less than half the number of sensors are compromised, bounds like (9) are very conservative in practice because the comparison functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ are hard to estimate.
To illustrate this point, we revisit Example 1 for the case where fewer than half the number of sensors are attacked.
Example 2.
Reconsider Example 1, but suppose sensor $1$ is under attack while sensors $2$ and $3$ are uncompromised. We have the number of sensors $p=3$ and the number of attacked sensors $q=1$, and, for the sake of this illustration, assume the attack signals $a_{2}(k)=a_{3}(k)=0$, for all $k\in\mathbb{Z}_{\geq 0}$. Furthermore, suppose the attack on sensor $1$ is $a_{1}(1)=-10.03$. We obtain the estimated set according to Algorithm 1 at time step $k=1$ as follows:
  • The measurement output set for each sensor is $\mathcal{Y}^{1}_{k}=[-6.1,-5.9]$ and $\mathcal{Y}^{3}_{k}=[3.83,4.03]$ as calculated before, and $\mathcal{Y}^{2}_{k}=[3.88,4.08]$ according to (5).
  • The time update set $\hat{\mathcal{X}}_{1|0}=[-20,20]$.
  • Measurement update sets according to (6): $\hat{\mathcal{Z}}^{1}_{1}=[-6.1,-5.9]$, $\hat{\mathcal{Z}}^{2}_{1}=[3.88,4.08]$, and $\hat{\mathcal{Z}}^{3}_{1}=[3.83,4.03]$.
  • Agreement sets using (7): for $\mathsf{J}_{1}=\{1,2\}$, $\mathsf{J}_{2}=\{1,3\}$ and $\mathsf{J}_{3}=\{2,3\}$, we get $\mathcal{I}^{1}_{1}=\hat{\mathcal{Z}}^{1}_{1}\cap\hat{\mathcal{Z}}^{2}_{1}=\emptyset$, $\mathcal{I}^{2}_{1}=\hat{\mathcal{Z}}^{1}_{1}\cap\hat{\mathcal{Z}}^{3}_{1}=\emptyset$, and $\mathcal{I}^{3}_{1}=\hat{\mathcal{Z}}^{2}_{1}\cap\hat{\mathcal{Z}}^{3}_{1}=[3.88,4.03]$.
  • Finally, the estimated set using (8) is $\hat{\mathcal{X}}_{1}=\mathcal{I}^{1}_{1}\cup\mathcal{I}^{2}_{1}\cup\mathcal{I}^{3}_{1}=[3.88,4.03]$.
We consider the center $3.955$ of $\hat{\mathcal{X}}_{1}$ as an estimate of $x(1)=4$, which gives the estimation error $0.045$. On the other hand, using the secure state estimation algorithm given in [12], the point-based estimate is obtained as $\hat{x}(1)=3.164$, which gives a much higher estimation error equal to $0.836$.
@@ 4.7 Bound on the Set-based Estimation Error
Since Theorem 2 guarantees that the true state $x(k)$ of system (3) lies in at least one of the agreement sets in the estimated collection of sets $\hat{\mathcal{X}}_{k}$ at each $k\in\mathbb{Z}_{\geq 0}$, it must also lie in a zonotope that overbounds $\hat{\mathcal{X}}_{k}$. That is, let $\hat{\mathcal{X}}_{k}$ be overbounded by a zonotope
@@
Then, considering $\hat{c}_{x}(k)$ to be an estimate of $x(k)$, the estimation error can be bounded by
@@
We remark that this error bound is significantly smaller in practice than the error bounds obtained by point-based secure estimators.
4.8 Strategies to Handle Computational Complexity
A major drawback of Algorithm 1 is that the attacker can design intelligent and stealthy attacks such that the number of agreement sets in the estimate $\hat{\mathcal{X}}_{k}$ keep growing with time. Although generating such attacks is a difficult problem as we argue in Section 5, the computational complexity of the algorithm may increase exponentially with time $k$ in the worst case scenario. Therefore, it is important to reduce the complexity by adopting several pruning methods that are described below.
An obvious step is to remove any empty or subsets of other sets in the estimated set $\hat{\mathcal{X}}_{k}$. It is also possible to obtain a single overbounding zonotope of $\hat{\mathcal{X}}_{k}$ as in (11), and use it instead of $\hat{\mathcal{X}}_{k}$ in the next time update. However, a better trade-off between accuracy and complexity is to not overbound the whole collection but only the intersecting zonotopes in the collection $\hat{\mathcal{X}}_{k}$. This may not make the cardinality of $\hat{\mathcal{X}}_{k}$ equal to one, but it reduces it significantly by allowing minimal loss of accuracy. Another method is to employ zonotope reduction methods [41] to reduce the number of generators in the zonotopes, which are often increased when performing the Minkowski sum operations.
@@ 5 Attack Detection, Identification, and Filtering
In this section, we provide sufficient conditions that guarantee the security of the S3E algorithm.
@@ 5.1 Attack Detection
The above proposition shows that whenever the estimated set contains more than one agreement sets that are disjoint, there has to be at least one sensor that is compromised. Once the attack is detected, the system operator can take necessary measures to secure the system and bring it to safety. If the attacker wants to stay stealthy, they must design attack signals that ensure that all the agreement sets yield a non-empty intersection. In the next subsections, we show that the magnitude of such a stealthy attack has to be sufficiently small.
@@ 5.2 Attack Identification
We say that the attack $a_{i}(k)$ is identified if the system operator knows for certain that sensor $i$ is compromised at time $k$, i.e., $i\notin\mathsf{S}_{k}$. The attack identification allows us to automatically discard the measurements coming from the compromised sensors in the S3E algorithm. We achieve attack identification through a measurement update (6), which involves a generalized intersection between the time update $\hat{\mathcal{X}}_{k|k-1}$ and the output measurement set $\mathcal{Y}_{k}^{i}$.
@@
Given a zonotope $C_{i}\bar{\mathcal{X}}_{k|k-1}$ with center $C_{i}c_{x}(k|k-1)$ and radius $\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$, we have that any point $p\in\mathbb{R}^{m_{i}}$ is outside the $C_{i}\bar{\mathcal{X}}_{k|k-1}$ if $\|C_{i}c_{x}(k|k-1)-p\|>\sqrt{m_{i}}\mathrm{rad}(C_{i}\bar{\mathcal{X}}_{k|k-1})$. Similarly, for $\mathcal{Y}_{k}^{i}$ whose center is $C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)$ and radius is $\mathrm{rad}(\mathcal{Y}_{k}^{i})$, any $p\in\mathbb{R}^{m_{i}}$ is outside $\mathcal{Y}_{k}^{i}$ if $\|C_{i}x(k)+v_{i}(k)-c_{v_{i}}+a_{i}(k)-p\|>\sqrt{m_{i}}\mathrm{rad}(\mathcal{Y}_{k}^{i})$. Here, the square root term $\sqrt{m_{i}}$ appears because the radius of a zonotope is a maximum norm $\|\cdot\|_{\infty}$, which relates to the Euclidean norm $\|\cdot\|$ as follows: for any vector $p\in\mathbb{R}^{m_{i}}$, $\|p\|\leq\sqrt{m_{i}}\|p\|_{\infty}$.
If the distance between the centers of $C_{i}\bar{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ is more than $\sqrt{m_{i}}[\mathrm{rad}(\bar{\mathcal{X}}_{k|k-1})+\mathrm{rad}(\mathcal{Y}_{k}^{i})]$, then any point $x\in\bar{\mathcal{X}}_{k|k-1}$ will be such that $C_{i}x\notin\mathcal{Y}_{k}^{i}$. Therefore, in this case, the measurement update (6) is empty. In other words, we have $\hat{\mathcal{Z}}_{k}^{i}=\emptyset$ if
@@
The sufficient condition in Theorem 4 provides a class of attack signals that can be identified at the measurement update step of the S3E algorithm. If the measurement update set $\hat{\mathcal{Z}}_{k}^{i}$ is empty, then it reveals that sensor $i$ is compromised at time $k$. However, the implication does not hold in the other direction. Nonetheless, while (13) establishes a threshold above which an attack will lead to an empty generalized intersection, it does not guarantee that smaller attacks will be undetected. The overapproximations used in the proof of Theorem 4 can introduce conservatism, meaning that attacks slightly below the threshold might still be detectable. In other words, the measurement update $\hat{\mathcal{Z}}_{k}^{i}$ could still be empty even if the attacker violates (13). This can happen if the attack signal is such that $C_{i}\hat{\mathcal{X}}_{k|k-1}$ and $\mathcal{Y}_{k}^{i}$ are inconsistent with each other at time $k$, resulting in an empty generalized intersection.
Notice that the bound (13) depends on the specific realization of measurement noise $v_{i}(k)$ at time $k$, which is unknown to both the attacker and the defender. One can further lower bound (13) and obtain a deterministic bound as follows:
@@
If the attacker wants to stay stealthy and effective, it can choose the attack signal $a_{i}(k)$, for every $i\in\mathbb{Z}_{[1,p]}$, satisfying
@@
Although this still does not ensure a non-empty $\hat{\mathcal{Z}}_{k}^{i}$, it may serve as a rough guideline for the attacker to generate a stealthy attack signal.